# Explain with Visual Key-Points Like a Real Mentor ! A Benchmark for Visual Solution-Explaining

### v1

@@ 5.2 Keypoint-based Explanation Generation
Math-PUMA 1.233 1.233 1.216
Gemini 2.0 Flash 3.749 3.343 3.992
@@
5.3 Does explanation quality reflect problem-solving ability?
For this analysis, ground-truth answers are withheld during inference. Models are first prompted to generate an explanation without access to the correct answer (“blind explanation”), and subsequently asked to produce a hypothetical answer based solely on their self-generated explanation.
Result
Table 5 (left) presents GPT-based evaluations (GPT-Eval) under the no-answer condition for the Keypoint-based Explanation Generation task, in which models are required to produce explanatory content without access to the final answer. Despite this constraint, performance patterns remain broadly consistent with those observed in previous experiments. However, the absolute scores drop across all models, indicating that the complexity of explanation generation hinges on the availability of the correct answer.
Table 5 (right) shows the accuracy of the answers that the models generated conditioned on their own explanations. Gemini 2.0 Flash, which previously demonstrated strong performance in visual solution-explaining, consistently outperforms other models across both Geometry and Graph tasks, achieving notably high accuracy in both MCQA and SA settings. In contrast, Math-PUMA performs poorly not only in Keypoint-based Explanation Generation but also in problem-solving accuracy. Notably, it is even outperformed by Qwen—a model not specifically optimized for mathematics—suggesting that the ability to generate coherent explanations may be linked to stronger problem-solving performance.
@@ 6 Conclusion
We introduce visual solution-explaining, a task that requires essential mathematics-teaching capabilities for delivering intuitive and insightful explanations alongside problem-solving. We also propose MathExplain, which assesses visual solution-explaining performance through two subtasks: (1) Visual Keypoint Identification, challenging models to generate visual keypoints that are crucial for subsequent explanations, and (2) Keypoint-based Explanation Generation, requiring models to integrate identified keypoints into coherent, solution-driven narratives. Our experimental results show that open-source models struggle to accurately identify the key visual elements necessary for effective problem comprehension. This shortcoming is consistently reflected in their explanatory text generation performance, particularly in mathematics-oriented models fine-tuned on domain-specific data. We believe that visual solution-explaining and MathExplain will inspire further research on developing models for educational applications, ultimately paving the way for real-world mentors capable of providing clear, student-friendly solutions. All code and datasets will be made publicly available.
@@ 8 Ethical Considerations
In this paper, we introduce the MathExplain benchmark, which is curated from an in-house mathematics education platform [^1] https://mathpresso.com/en. During the data collection process, all materials were carefully curated to ensure full compliance with copyright regulations. For data annotation, we employed individuals who are proficient in English and hold undergraduate degrees, ensuring that they possess adequate mathematical problem-solving skills.
@@ Appendix A Experiments Details
In this study, experiments were conducted using the LMMs-eval repository (Li et al., 2024). This repository provides a comprehensive framework for evaluating multi-modal models across various tasks. In all GPT evaluations, we used the gpt-4o-2024-08-06 endpoint. When conducting experiments with open-source multimodal models, we leveraged the official implementation codes in conjunction with publicly available weights from the Huggingface Hub  [^2] https://huggingface.co/models. The following model parameters were used for each model:
@@
These models were evaluated in our benchmark, which included tasks designed to assess both visual understanding and textual explanation capabilities. The selection of models spans a range of architectures and performance levels, providing insights into current advancements in multi-modal learning.
@@ A.1 Computational Resources
For closed-source models, we used the OpenAI API and Gemini Developer API to infer the output of GPT-4o and Gemini 2.0. For open-source models such as Math-PUMA, URSA, LLaVA, QWEN-2.5-VL, and Molmo, inference was performed using a NVIDIA A6000 48GB GPU. While the exact inference speed varies depending on the task, model, and lengths of prompts and responses, a query takes about 50 seconds to be answered.
A.2 Evaluation on automatic metrics
@@
Table 6: Automated Evaluation Metrics for the Generating Explanation Text Task on MathExplain

### v2

@@ 5.3 Does explanation quality reflect problem-solving ability?
Math-PUMA 1.268 1.094 1.058
Gemini 2.0 Flash 3.598 3.292 3.867
@@
Table 6 (right) shows the accuracy of the answers that the models generated conditioned on their own explanations. Gemini 2.0 Flash, which previously demonstrated strong performance in visual solution-explaining, consistently outperforms other models across both Geometry and Graph, achieving notably high accuracy in both MCQA and SA settings. In contrast, Math-PUMA performs poorly not only in Keypoint-based Explanation Generation but also in problem-solving accuracy. Notably, it is even outperformed by Qwen, a model not specifically optimized for mathematics, suggesting that the ability to generate coherent explanations may be linked to stronger problem-solving performance.
@@ 6 Conclusion
We introduce visual solution-explaining task, that requires essential mathematics-teaching capabilities for delivering intuitive and insightful explanations alongside problem-solving. We also propose MathExplain, which assesses visual solution-explaining performance through two subtasks: (1) Visual Keypoint Identification, challenging models to generate visual keypoints that are crucial for subsequent explanations, and (2) Keypoint-based Explanation Generation, requiring models to integrate identified keypoints into coherent, solution-driven narratives. Our experimental results show that open-source models struggle to accurately identify the key visual elements necessary for effective problem comprehension. This shortcoming is consistently reflected in their explanatory text generation performance, particularly in mathematics-oriented models. We believe that visual solution-explaining and MathExplain will inspire further research on developing models for educational applications, ultimately paving the way for real-world mentors capable of providing clear, student-friendly solutions. All code and datasets will be made publicly available.
@@ 8 Ethical Considerations
In this paper, we introduce the MathExplain benchmark, curated from an in-house mathematics education platform [^1] https://mathpresso.com/en. During the data collection process, all materials were carefully reviewed to ensure full compliance with copyright regulations. For the data annotation, we employed annotators who are proficient in English and hold undergraduate degrees, thereby ensuring that they possess the necessary mathematical problem-solving skills.
@@ Appendix A Experiments Details
In this study, experiments were conducted using the LMMs-eval repository (Li et al., 2024). This repository provides a comprehensive framework for evaluating multimodal models across various tasks. In all GPT evaluations, we used the gpt-4o-2024-08-06 endpoint. When conducting experiments with open-source multimodal models, we leveraged the official implementation codes in conjunction with publicly available weights from the Huggingface Hub  [^2] https://huggingface.co/models. The following model parameters were used for each model:
@@
These models were evaluated in our benchmark, which included tasks designed to assess both visual understanding and textual explanation capabilities. The selection of models spans a range of architectures and performance levels, providing insights into current advancements in multimodal learning.
@@ A.1 Computational Resources
For closed-source models, we used the OpenAI API and Gemini Developer API to infer the output of GPT-4o and Gemini 2.0 Flash. For open-source models such as Math-PUMA, URSA, LLaVA, QWEN-2.5-VL, and Molmo, inference was performed using a NVIDIA A6000 48GB GPU. While the exact inference speed varies depending on the task, model, and lengths of prompts and responses, a query takes about 50 seconds to be answered.
A.2 Human Evaluation Setup
Three evaluators, all holding a bachelor’s or master’s degree in engineering, are assessing AI model outputs for 80 problems. Specifically, they are evaluating the results produced by the Math-PUMA, Qwen2.5-VL, and Gemini 2.0 Flash models, which represent the OM, OG, and CG categories. Each criterion is being rated on a five-point Likert scale.
A.3 Evaluation on Conventional Text Similarity Metrics
@@
Table 7: Automated Evaluation Metrics for the Keypoint-based Explanation Generation Task on MathExplain.